- Press Release
- Dec 2, 2022
NASA Space Station On-Orbit Status 8 February 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
Cosmonaut Sharipov conducted the fifth and final experiment of the current (12th) science session with the Russian/German TEKh-20 Plasma Crystal-3 (PK-3) payload, modified from yesterday’s run. [After wakeup, Salizhan again activated the PK-3/N turbopump in the Service Module (SM) Transfer Compartment (PkhO), to keep the work chamber (ZB) in the SM RO (Work Compartment) at a vacuum. Then he set up the experiment, supported by tagup with the ground via S-band. Video recording began ~20 min after experiment initiation. The experiment, running in automatic mode, is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber. Main objective today was to obtain plasma dust clouds at various pressures and particle quantity with and without superimposition of an LF (low-frequency) harmonic electrical field at variable frequency. Disassembly of PK-3 was to be performed from today’s discretionary task list. Note: As Sharipov stated, Cosmonauts appreciate PK-3 particularly because it allows them to actually see the results of their work, as opposed to past crystal growth experiments.]
At ~1:40am EST, CDR/SO Leroy Chiao began the “bakeout” regeneration of two expended EMU Metox (metal oxide) recyclable CO2 absorption canisters in the “Quest” Airlock (A/L). As part of his Metox ops, Chiao collected two data sets of temperatures in the A/L during specific phases of the regen cycle, both times taking 10-second readings at 12 different locations using the Velocicalc instrument. Metox regen ends at ~3:50pm. [Since the regeneration process releases CO2 (carbon dioxide) in the cabin, the CDRA (CO2 removal assembly) was activated by the ground overnight, as well as the A/L CCAA (common cabin air assembly) air conditioner. In the past, the ground had set the CCAA at a low setpoint of 65 degF (18.3 degC) for the high temperature expected from the oven during the “baking” phase, but crewmembers found the A/L uncomfortably cold during this activity. Today’s readings will help the ground to determine whether the cabin set point can be relaxed in the future. The Metox regeneration provides two fresh canisters for LF-1 (STS-114), which has three EVAs scheduled.]
Later, Leroy conducted the weekly data take with the two 16P-delivered CSA-CP units (Compound Specific Analyzer-Combustion Products) #1018 & #1019, with the purpose to monitor the desired ongoing decontamination (outgassing) of the deployed instruments. [CSA-CP measures O2 (oxygen), CO (carbon monoxide), HCN (hydrogen cyanide), and HCl (hydrogen chloride).]
Continuing the current outfitting to modify and prepare the Russian ASN-M satellite navigation system in the SM for the European ATV (Automated Transfer Vehicle), continuing through 2/11, the FE today routed and connected the onboard cable network (BKS) to the four NPM receiver module units installed yesterday. The work was again supported by S-band tagup with ground specialists. [In tomorrow’s ASN activity, Salizhan will install two navigation computing units (NVM-1,2), followed on 2/10 by connecting the ASN-M to the SUBA onboard equipment control system and on 2/11 to the SBI onboard measuring system, completing the installation. Although not yet operational, ASN-M uses GLONASS satellites (the Russian GPS equivalent) to update the ISS state vector (SV, position & velocity plus time) without using the ground (which currently has to uplink daily SV updates) or requiring SV transfers from the U.S. segment (USOS) from time to time. The ASN equipment was originally factory-installed in the SM but was found faulty and had to be returned to the ground. After repair it was shipped again to the station on Progress 11 and re-installed by Yuri Malenchenko on 7/8/03, followed by various troubleshooting attempts and the current ATV mods.]
On the MedOps defibrillator equipment, successfully checked out yesterday, Leroy initiated the regular recharge process on battery #1 to full capacity and subsequently performed it also on battery #2. The task concluded with a battery voltage check. [Each battery was charged for ~3.5 hrs, and its open-circuit voltage was tested at the end with the Aeolus volt/amp scopemeter, then removed and stowed again. Nominally, the defib has a battery installed at all times, but with this particular unit the PDIM (power data interface module) is not functioning properly and would overcharge the batteries if left inside. They have to be charged every 60 days along with the defib checkout.]
After the Science Officer set up the ADUM (Advanced Diagnostic Ultrasound in Microgravity) experiment and performed some preparatory HRF (Human Research Facility) laptop file management, Salizhan used the PC and Ethernet/OpsLAN for an ADUM OPE (On-board Proficiency Enhancer) session in preparation for another round of bone (Scan Z) and abdominal (Scan B) scans on himself, scheduled for 2/11 (Friday). [The FE used the ADUM OPE compact disk on the PC, focusing on cardiac, thoracic & bone scanning, plus data acquisition (probe positioning) and principles of remote guidance, ultrasound, and anatomy.]
The SO also conducted another monthly photography session (Session 9) of the BCAT-3 (Binary Colloid Alloy Test 3) science investigation. Afterwards, the equipment was dismantled again. [After setting up the SGSM (slow growth sample module) in the Lab MWA (Maintenance Work Area) at EXPRESS Rack 5 (ER5), Chiao photographed Group B samples (4, 5) and Group C (8, 9, and 10). The ground will review the images, comparing them with earth-based crystallization samples running in parallel.]
A new list of “Saturday Science” options for Dr. Chiao was uplinked for his selection later tonight. [The options for 2/12 are HRF FOOT (Foot/Ground Reaction Forces during Space Flight) experiment kit transfers, FMVM (Fluid Merging Viscosity Measurement) experiment ops, and CBOSS-FDI (Cellular Biotechnology Support Systems-Fluid Dynamics Investigation) Tissue Culture Module (TCM) and syringe bubble removal, with images.]
Leroy completed the remaining portion of the current inventory/audit of the crew’s computer accessories “pantry”. Afterwards he tagged up with ground specialists via S-band to discuss the results.
Chiao performed his daily check of the Total Dose reading and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference). [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPC’s temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]
The CDR conducted the routine SOZh/ECLSS servicing/inspection in the SM, including ASU toilet facilities maintenance, and, working off the “job jar” task list, prepared the regular IMS “delta” file for the daily automated export/import to the three IMS databases on the ground.
The crew worked out in their daily 2.5-hr. physical exercise program on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (both aerobic and anaerobic). Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 1 of a new set) and one hour on VELO.
Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
Leroy also completed the weekly TVIS maintenance, which generally checks the condition of the SPDs (subject positioning devices) and records time & date. [For running on the treadmill (motor-powered or passive), the crewmember wears a special harness with bungees that are hooked into the strut-like SPDs, one left, one right, to keep him centered and minimize the force transferred to the station during exercise, while keeping his feet in contact with the running surface.]
Working off the Russian task list, Salizhan transferred the first service data of the “RokvISS” payload from the BSPN payload server, along with BSPN log files and BRI (RS OpsLAN/Ethernet SmartSwitch) control software files via the ISS Wiener laptop to a PCMCIA flash card for subsequent downlink on OCA comm.
At ~5:20am, the crew conducted a 20-min. interactive TV conference with Anatoly Perminov, Director of Roskosmos, and Edelgard Bulmahn, German Minister of Education and Science, at TsUP/Moscow, on K-band/video and S-band/audio relayed via MCC-H. [Subjects of discussion were the Plasma Crystal and RokvISS experiments being conducted on the Russian segment with Germany’s participation. Sharipov pointed out the PK-3’s “Telescience” computer video control monitor (VKU) and vacuum work chamber (ZV). The German-developed high-tech robotic arm RokvISS is currently being tested from its ground control center at Waldheim (suburb of Munich), which is reportedly receiving a steady signal and the first images from the hardware video camera. Transition from the tests to the nominal ops phase of RokvISS is planned for March ’05.]
At ~10:40am, the crew also engaged in a educational/PAO event, exchanging Q&A’s with grade 4-6 students at Crossroads Elementary School, St. Paul, MN, with a number of Minnesota VIPs present. A webcast was also planned.
At ~10:40am EST, ISS attitude control was handed over to the Russian SUD motion control system (MCS), followed by a thruster-effected maneuver to XPOP TEA (x-axis perpendicular to orbit plane/torque equilibrium attitude) at 10:47am, followed by return of control authority to the U.S. CMGs (control moment gyroscopes) at ~11:07am. Autotrack bias setting of the P6 solar array wings (for drag reduction) was subsequently (12:30pm) configured to 37.5 deg for BGA (Beta gimbal assembly) 2B and -37.5 deg for BGA 4B.
Today’s CEO (crew earth observations) photo targets, still for LVLH attitude, were Berlin, Germany (unusual winter clearing was forecast. Looking a touch left for the city center. Berlin is large but difficult to identify due to its numerous large forested parks), Lake Nasser, Toshka Lakes, Egypt (a detailed mapping swath of these lakes in southwest Egypt was requested. All the lakes except the one nearest the Nile/Lake Nasser were unintended extensions of the first lake. The lakes are intended to provide irrigation for ~20 million Egyptians from the overpopulated Nile delta and valley), and Addis Ababa, Ethiopia (looking right for an oblique view of this seldom seen highland city).
CEO photography can be viewed and studied at the websites:
See also the website “Space Station Challenge” at:
To view the latest photos taken by the expedition 10 crew visit:
- http://spaceflight.nasa.gov/gallery/images/station/crew-10/ndxpage1.html at NASA’s Human Spaceflight website.
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
- ISS Reboost — 2/15 (~8:22pm EST, ~1.8 m/s; phasing for 17P launch);
- Progress M-51 (16P) undocking & destructive reentry — 2/26/05;
- Progress M-52 (17P) launch — 2/28/05.
- EVA-13 — 3/25/05;
- Soyuz TMA-6 (10S) launch — 4/15/05 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
- Soyuz TMA-5 (9S) undock — 4/25/05 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
- Progress M-53 (18P) launch — 6/10/05;
- Progress M-54 (19P) launch — 8/24/05;
- Soyuz TMA-7 (11S) launch — 9/27/05.
ISS Altitude History
Apogee height — Mean Altitude — Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.